# Stability of oppositely-propagating pair of Hill’s spherical vortices

### v1

@@ Abstract.
We establish the stability of a pair of Hill’s spherical vortices moving away from each other in 3D incompressible axisymmetric Euler equations without swirl. Each vortex in the pair propagates away from its odd-symmetric counterpart, while keeping its vortex profile close to Hill’s vortex. This is achieved by analyzing the evolution of the interaction energy of the pair and combining it with the compactness of energy-maximizing sequences in the variational problem concerning Hill’s vortex. The key strategy is to confirm that, if the interaction energy is initially small enough, the kinetic energy of each vortex in the pair remains so close to that of a single Hill’s vortex for all time that each vortex profile stays close to the energy maximizer: Hill’s vortex. An estimate of the propagating speed of each vortex in the pair is also obtained by tracking the center of mass of each vortex. The estimate can be understood as optimal in the sense of the power exponent of the $\varepsilon$—the small perturbation measured in the ($L^{1}\cap L^{2}$+impulse) norm—appearing in the error bound cannot be improved.
@@ 1. Introduction
in cylinderical coordinates ${\bf x}=(r,\theta,z)$. Then we can reduce (1.1) to the following active scalar equation
@@ 1.1. Main statement
Theorem 1.2 below contains the shift function estimate. That is, with some additional assumptions on $\xi_{0}$, we establishes that any shift function $\tau:[0,\infty)\to[1,\infty)$ in Theorem 1.1 propagates approximately with the speed $W_{H}=2/15$, which is the traveling speed of Hill’s spherical vortex $\xi_{H}$, i.e.,
@@
The shift estimate (1.3) in Theorem 1.2 above is sharp, in the sense of the power exponent of $\varepsilon$ (which is 111) in the right-hand. Indeed, when considering the odd-pair of $\lambda\xi_{H}$ ($\lambda$-strength Hills vortex) as a perturbed data in Theorem 1.1, its $L^{1}\cap L^{2}\cap L^{1}_{w}$ error ($\sim|\lambda-1|$) in Theorem 1.1 coincides with the shift error ($\sim|\lambda-1|(t+1)$), knowing that the traveling speed of $\lambda\xi_{H}$ is given as $\lambda W_{H}$. We note that the method used to obtaining the estimate (1.3) can be used to improve the shift function estimate given in [20, Proposition 3.2], where the authors considered a single Hill’s vortex, in the optimal way of raising the power exponent of $\varepsilon$ from 1/21/21 / 2 to 111.
@@
Thanks to the shift function estimate (1.3) together with the stability given in Theorem 1.1, one can prove the linear-in-time filamentation near the odd-symmetric pair of the Hill’s vortices moving away from each other nearly with speed $W_{H}=2/15$. Indeed, we refer to [20] for the result of filamentation near a single Hill’s vortex. It was obtained from the stability of a single Hill’s vortex (given in [17]) and an additional shift function estimate. For the similar filamentation result near the single Chapyglin-Lamb’s dipole in $\mathbb{R}^{2}$, we refer to [19], and see also [23] for the filamentation result near a collection of the Sadovskii vortex patches.
@@
with the cylinderical coordinate $(r,\theta,z)$ where the unit vectors are defined by
@@
we say the velocity is axisymmetric (about $z$-axis) when each components $u^{r},u^{\theta},u^{z}$ does not depend on $\theta$. The no-swirl condition stands for $u^{\theta}\equiv 0$. In this setting, the velocity ${\bf u}$ has the form
@@
Then the 3D Euler equations (1.1) is reduced to (1.2):
@@ 1.3.1. Interaction between two vortices
For the governing equation (1.2), the kinetic energy $E[\xi]$ is conserved in time $t\geq 0$, and has a bi-linear form in $\xi$:
@@ 1.3.2. Shift function estimate
Since $\xi_{H}$ travels in time with the constant velocity $W_{H}{\bf e}_{z}$ with $W_{H}=2/15$, we expect that the shift fucntion $\tau(t)$ is close enough to $W_{H}t$ in a finite time interval (Theorem 1.2), i.e.,
@@ 1.4.3. Multi-vortex problems
The key strategy is to construct a configuration of well‑separated vortex objects and prove that the residual interaction is too small to destroy their rigidity/coherency. In vortex dynamics, canonical building blocks might include translating dipoles (e.g., Chaplygin-Lamb dipole, Sadovskii vortex patches), rotating V-states (e.g., Rankine vortex, Kirchhoff ellipse), axisymmetric vortex rings including Hill’s vortex, and desingularized point vortex (or point-vortex ring). The question is whether copies of such structures can coexist, either moving apart or forming choreographed motions, without significant deformation. Several recent works has used the Chaplygin-Lamb dipole as a building block to construct and solve multi-vortex problems; see [22] for the case of the dipole pair moving away from each other, and see [2] for the stability of finite sums of the dipoles moving together in one direction, in which the faster dipoles are positioned to the right of the slower ones.
@@ 1.5.3. Other types of multi-vortex
Even though the current paper concerns the pair of Hill’s vortex, it can be generalized to other types of vortices that are characterized as energy maximizers under certain constraints; e.g., see [22, Remark 2.9] for a similar discussion. One can also try to increase the number of vortices can be more than two, provided that the interaction decreases. As far as the interaction energy vanishes, the vortices will nearly retain their initial shapes, implying their stability.
@@ 2.2. Anti-parallel flow
In subsection 1.2, the 3D axisymmetric Euler equations with no swirl has the reduced form (1.2):
@@
In other words, we consider the anti-parallel flow that is non-negative in the upper-half space +$\mathbb{R}^{3}_{+}$. In this case, we identify the non-negative part $\langle\xi\rangle_{+}$ of the solution $\xi$ as the restriction of $\xi$ to the half space +$\mathbb{R}^{3}_{+}$. Then the following lemma shows that the impulse of such a solution decreases strictly in time (We do not claim any novelty of this result. For example, it can be found from [18, Theorem 1.1]. Still, we give the proof for completeness).
@@ 2.3. Variational problem concerning Hill’s spherical vortex
In this section, we give some results in [17] about the variational problem concerning the Hill’s spherical vortex $\xi_{H}$. We first introduce a scaling family of Hill’s vortex.
@@
By the scaling invariance of Euler equations, a family of traveling solutions of two parameters $\lambda,a>0$ is obtained from the scaling
@@ 2.4. Compactness of the set of maximizers
the unique weak solution $\xi(t)$ of (1.2) satisfies
@@
Without loss of generosity, from Theorem 2.5 we put $\xi=\xi_{H}^{\mu}$ to write
@@
Let $\varepsilon>0$ be any constant. We fix a small $\mu:=\mu_{0}/2\in(0,\mu_{0})$ for the constant $\mu_{0}>0$ given in Theorem 3.1 to guarantee $\kappa_{\mu}<1$ (if necessary, we redefine $\mu_{0}$ smaller enough). We define $\delta,d_{0}>0$ as
@@
Here, we define a bi-liner operator $\mathcal{B}$ defined as
@@
and the fact that $\partial_{r}G(r,r^{\prime},z,z^{\prime})$ is a function of triple of variables $(r,r^{\prime},|z-z^{\prime}|)$, the bi-liner operator $\mathcal{B}$ has the translation invariance in $z$-axis in that, by denoting the shift operator as $f^{\tau}:=f(\cdot-\tau{\bf e}_{z})$ for $\tau\in\mathbb{R}$, we have
@@
Then, using (i)-(iii) above, we will obtain (4.12). First, we combine (4.14) and (4.16) to put we obtain

### v2

@@ Abstract.
We establish the stability of a pair of Hill’s spherical vortices moving away from each other in 3D incompressible axisymmetric Euler equations without swirl. Each vortex in the pair propagates away from its odd-symmetric counterpart, while keeping its vortex profile close to Hill’s vortex. This is achieved by analyzing the evolution of the interaction energy of the pair and combining it with the compactness of energy-maximizing sequences in the variational problem concerning Hill’s vortex. The key strategy is to confirm that, if the interaction energy is initially small enough, the kinetic energy of each vortex in the pair remains so close to that of a single Hill’s vortex for all time that each vortex profile stays close to the energy maximizer: Hill’s vortex. An estimate of the propagating speed of each vortex in the pair is also obtained by tracking the center of mass of each vortex. The estimate can be understood as optimal in the sense that the power exponent of the $\varepsilon$—the small perturbation measured in the ($L^{1}\cap L^{2}$+impulse) norm—appearing in the error bound cannot be improved.
@@ 1. Introduction
in cylindrical coordinates ${\bf x}=(r,\theta,z)$. Then we can reduce (1.1) to the following active scalar equation
@@ 1.1. Main statement
Theorem 1.2 below contains the shift function estimate. That is, with some additional assumptions on $\xi_{0}$, we establish that any shift function $\tau:[0,\infty)\to[1,\infty)$ in Theorem 1.1 propagates approximately with the speed $W_{H}=2/15$, which is the traveling speed of Hill’s spherical vortex $\xi_{H}$, i.e.,
@@
The shift estimate (1.3) in Theorem 1.2 above is sharp, in the sense of the power exponent of $\varepsilon$ (which is 111) in the right-hand side. Indeed, when considering the odd-pair of $\lambda\xi_{H}$ ($\lambda$-strength Hills vortex) as a perturbed data in Theorem 1.1, its $L^{1}\cap L^{2}\cap L^{1}_{w}$ error ($\sim|\lambda-1|$) in Theorem 1.1 coincides with the shift error ($\sim|\lambda-1|(t+1)$), knowing that the traveling speed of $\lambda\xi_{H}$ is given as $\lambda W_{H}$. We note that the method used to obtaining the estimate (1.3) can be used to improve the shift function estimate given in [20, Proposition 3.2], where the authors considered a single Hill’s vortex, in the optimal way of raising the power exponent of $\varepsilon$ from 1/21/21 / 2 to 111.
@@
Thanks to the shift function estimate (1.3) together with the stability given in Theorem 1.1, one can prove the linear-in-time filamentation near the odd-symmetric pair of the Hill’s vortices moving away from each other nearly with speed $W_{H}=2/15$. Indeed, we refer to [20] for the result of filamentation near a single Hill’s vortex. It was obtained from the stability of a single Hill’s vortex (given in [17]) and an additional shift function estimate. For the similar filamentation result near the single Chaplygin-Lamb’s dipole in $\mathbb{R}^{2}$, we refer to [19], and see also [23] for the filamentation result near a collection of the Sadovskii vortex patches.
@@
with the cylindrical coordinate $(r,\theta,z)$ where the unit vectors are defined by
@@
we say the velocity is axisymmetric (about $z$-axis) when each component $u^{r},u^{\theta},u^{z}$ does not depend on $\theta$. The no-swirl condition stands for $u^{\theta}\equiv 0$. In this setting, the velocity ${\bf u}$ has the form
@@
Then the 3D Euler equations (1.1) are reduced to (1.2):
@@ 1.3.1. Interaction between two vortices
For the governing equation (1.2), the kinetic energy $E[\xi]$ is conserved in time $t\geq 0$, and has a bilinear form in $\xi$:
@@ 1.3.2. Shift function estimate
Since $\xi_{H}$ travels in time with the constant velocity $W_{H}{\bf e}_{z}$ with $W_{H}=2/15$, we expect that the shift function $\tau(t)$ is close enough to $W_{H}t$ in a finite time interval (Theorem 1.2), i.e.,
@@ 1.4.3. Multi-vortex problems
The key strategy is to construct a configuration of well‑separated vortex objects and prove that the residual interaction is too small to destroy their rigidity/coherency. In vortex dynamics, canonical building blocks might include translating dipoles (e.g., Chaplygin-Lamb dipole, Sadovskii vortex patches), rotating V-states (e.g., Rankine vortex, Kirchhoff ellipse), axisymmetric vortex rings including Hill’s vortex, and desingularized point vortex (or point-vortex ring). The question is whether copies of such structures can coexist, either moving apart or forming choreographed motions, without significant deformation. Several recent works have used the Chaplygin-Lamb dipole as a building block to construct and solve multi-vortex problems; see [22] for the case of the dipole pair moving away from each other, and see [2] for the stability of finite sums of the dipoles moving together in one direction, in which the faster dipoles are positioned to the right of the slower ones.
@@ 1.5.3. Other types of multi-vortex
Even though the current paper concerns the pair of Hill’s vortex, it can be generalized to other types of vortices that are characterized as energy maximizers under certain constraints; e.g., see [22, Remark 2.9] for a similar discussion. One can also try to increase the number of vortices to more than two, provided that the interaction decreases under certain configurations. As far as the interaction energy vanishes, the vortices will nearly retain their initial shapes, implying their stability.
@@ 2.2. Anti-parallel flow
In subsection 1.2, the 3D axisymmetric Euler equations with no swirl have the reduced form (1.2):
@@
In other words, we consider the anti-parallel flow that is non-negative in the upper half-space +$\mathbb{R}^{3}_{+}$. In this case, we identify the non-negative part $\langle\xi\rangle_{+}$ of the solution $\xi$ as the restriction of $\xi$ to the half space +$\mathbb{R}^{3}_{+}$. Then the following lemma shows that the impulse of such a solution decreases strictly in time (We do not claim any novelty of this result. For example, it can be found from [18, Theorem 1.1]. Still, we give the proof for completeness).
@@ 2.3. Variational problem concerning Hill’s spherical vortex
In this section, we give some results in [17] about the variational problem concerning Hill’s spherical vortex $\xi_{H}$. We first introduce a scaling family of Hill’s vortex.
@@
By the scaling invariance of Euler equations, a family of traveling solutions parametrized by $\lambda,a>0$ is obtained from the scaling
@@ 2.4. Compactness of the set of maximizers
the axisymmetric weak solution $\xi(t)$ of (1.2) satisfies
@@
Without loss of generality, from Theorem 2.5 we put $\xi=\xi_{H}^{\mu}$ to write
@@
Let $\varepsilon>0$ be any constant. We fix a small $\mu:=\mu_{0}/2\in(0,\mu_{0})$ for the constant $\mu_{0}>0$ given in Theorem 3.1 to guarantee $\kappa_{\mu}<1$ (if necessary, we redefine $\mu_{0}$ small enough). We define $\delta,d_{0}>0$ as
@@
Here, we define a bilinear operator $\mathcal{B}$ as
@@
and the fact that $\partial_{r}G(r,r^{\prime},z,z^{\prime})$ is a function of triple of variables $(r,r^{\prime},|z-z^{\prime}|)$, the bilinear operator $\mathcal{B}$ has the translation invariance in $z$-axis in that, by denoting the shift operator as $f^{\tau}:=f(\cdot-\tau{\bf e}_{z})$ for $\tau\in\mathbb{R}$, we have
@@
Then, using (i)-(iii) above, we will obtain (4.12). First, we combine (4.14) and (4.16) to observe that